OpenStudy (anonymous):

Which ordered pair is on the graph of the equation 3x + 7y = 11? A. (0,–4) B. (0, 4) C. (6, –1) D. (1, –2)

OpenStudy (solomonzelman):

plug each of the coordinates in, I'll show you on choice A. 3x + 7y = 11 (0,–4) 3(0)+7(-4)=11 -28=11 That can't be so A doesn't work, PLUG IN the REST.

OpenStudy (anonymous):

so its b

OpenStudy (solomonzelman):

No, it's not b.

OpenStudy (anonymous):

its not b

OpenStudy (solomonzelman):

3(0)+7(4)=11 28=11 See why it's not b? Plug in C and D, see which of them works.

OpenStudy (anonymous):

are you gonna try ??

OpenStudy (anonymous):

its c

OpenStudy (solomonzelman):

Yup

OpenStudy (anonymous):

thanks

OpenStudy (solomonzelman):

Anytime!